# Discovery of Probabilistic Dirichlet-to-Neumann Maps on Graphs

### v1

@@ 1 Introduction
Multiphysics simulations are often composed of independent submodules exchanging information or fluxes through shared interfacial states [12]. A modular “systems-of-systems” perspective is therefore attractive for the prediction, design, and control of large-scale systems whose sheer complexity renders traditional simulation intractable. For example, the Energy Exascale Earth System Model (E3SM) forecasts planetary climate dynamics by coupling submodels representing the atmosphere, ocean, sea and land ice, and terrestrial systems. These submodels interact by exchanging mechanical and transport fluxes determined by their physical states [24]. Similarly, embedded electronic systems rely on individual circuit components that exchange voltage and current information, enabling the construction of integrated systems made up of billions of transistors.
@@
This exchange of state across subsystem boundaries can be formalized using Dirichlet-to-Neumann (D2N) maps. These maps arise in linear elliptic boundary value problems, where solving a partial differential equation (PDE) with given Dirichlet boundary conditions yields corresponding fluxes at those boundaries [27]. D2N maps thus provide a natural abstraction for characterizing interactions between submodels sharing common interfaces. Many established domain decomposition techniques—including mortar methods [1, 4], FETI methods [13], and hybridizable methods [8]—can be viewed within the D2N framework. The rigorous theoretical foundations underpinning traditional D2N maps offer convergence and stability guarantees when coupling large legacy simulation codes [19, 34]. However, the evaluation of a D2N map traditionally requires computationally intensive solves of the governing equations within each subdomain, which quickly becomes prohibitively expensive in multi-physics and multiscale settings.
@@
Recent efforts have therefore turned to data-driven schemes that replace costly subdomain models with a learned approximation of the D2N map [5, 15, 19]. While such surrogates dramatically reduce computational cost, they generally lack the theoretical guarantees of traditional D2N methods. The absence of rigorous a-priori error or stability guarantees for data-driven D2N surrogates poses significant challenges for verification, validation, and reliable deployment of learned models.
@@
We address this gap by combining discrete exterior calculus (DEC) [37] with computational graph completion (CGC) — a generalization of the optimal recovery problem (ORP) to the nonlinear graph setting [28]. Our framework learns flux-state relationships on graphs while producing analytically and computationally tractable posterior distributions that characterize uncertainty in the D2N map. Together, CGC and DEC deliver an efficient optimization strategy along with closed-form expressions for prediction error and out-of-distribution inference, re-introducing rigorous guarantees into data-driven D2N coupling.
@@ 2 Optimal recovery on graphs via Gaussian processes
[…]
@@
This topic is well-explored in the literature [11, 25, 26, 29]. Variations of this problem include sensor reconstruction problems (e.g., [23]) and boundary value inverse problems (e.g., [21]). More recently, the computational graph completion (CGC) framework [28] used Gaussian processes (GPs) to generalize the ORP to the graph setting, where constraints may be nonlinear.
GPs offer a powerful probabilistic framework for approaching the optimal recovery problem. A GP defines a prior distribution over a space of functions that can be updated with observed data to obtain a posterior distribution. The posterior mean provides the optimal recovery of the unknown function given the data, while the posterior covariance quantifies uncertainty. This makes GPs particularly useful for applications where both function approximation and uncertainty quantification are crucial. In this section, we provide a brief overview of Gaussian process regression, but we refer the interested reader to [31] for a more comprehensive but accessible treatment.
@@
$\begin{bmatrix}f(\textbf{x})\\ \mathbf{Y}\end{bmatrix}\sim\mathcal{N}\left(\mathbf{0},\begin{bmatrix}K(\textbf{x},\textbf{x})&K(\textbf{x},\textbf{X})\\ K(\textbf{X},\textbf{x})&K(\textbf{X},\textbf{X})\end{bmatrix}\right).$ (5)
@@
A GP is characterized by a covariance (kernel) function $K(\cdot,\cdot)$, which must be symmetric and positive-definite. According to the Moore–Aronszajn Theorem, each kernel $K(\cdot,\cdot)$ uniquely induces a corresponding reproducing kernel Hilbert space (RKHS) $\mathcal{H}_{K}$ [2]. This RKHS contains precisely the functions spanned by the GP kernel, and functions drawn from the GP prior belong to $\mathcal{H}_{K}$ with high probability. Working in $\mathcal{H}_{K}$ therefore yields solutions that are regularized according to the smoothness properties inherited from $K(\cdot,\cdot)$. Thus, prior information about the nature of the true function $f$ can be incorporated into the choice of kernel. The squared exponential RBF kernel:
@@
in the presence of noise. By the Representer Theorem, the solution to this problem lies in the span of kernel evaluations at the data points, dramatically simplifying the problem [33]. Indeed, the GP posterior mean is exactly the minimizer of the RKHS-regularized problem, ensuring both smoothness and consistency with the data. In 1, we state the theorem that yields a closed-form solution for this minimization problem.
@@ 3 Discrete exterior calculus
DEC extends the framework of differential geometry and exterior calculus to discrete structures like graphs, providing a precise language through which we can define differential operators (e.g., gradient, divergence, and curl) in a discrete setting [10, 18]. It is particularly well-suited for modeling flows on graphs, where quantities such as mass, heat, or electrical charge must obey conservation laws. In prior work, DEC has been used to learn physics from data [37]. In our framework, we leverage DEC to constrain our graph-based GP recovery problem, ensuring that the learned functions obey conservation principles across the graphs we consider. This integration yields physically meaningful predictions while maintaining the flexibility of GPs.
@@
$(\mathbf{D}_{0})_{e,v}=\begin{cases}-1&\text{if $v$ is the source of edge $e$},\\ 1&\text{if $v$ is the target of edge $e$},\\ 0&\text{otherwise}.\end{cases}\$ (13)
@@ 4.2 Formulation of the optimization problem
$\begin{split}\min_{\theta,\mathbf{u}_{\text{un}}}\min_{\mathbf{F}_{\text{un}}}\sum_{e\in\mathcal{E}}&\min_{f_{e}\in\mathcal{H}_{K_{e}}}\left\lVert f_{e}\right\rVert_{K_{e}}^{2}+\frac{1}{\sigma_{\epsilon}^{2}}\left\lVert f_{e}(\mathbf{u}_{e})-\mathbf{F}_{e}\right\rVert_{2}^{2}+\log\det(K_{e}(\mathbf{u}_{e},\mathbf{u}_{e})+\sigma_{\epsilon}^{2}I)\\ &\text{s.t.}\qquad\delta_{0}^{\intercal}\mathbf{F}=0,\\ \end{split}$ (21)
@@
$\begin{split}\min_{\theta,\mathbf{u}_{\text{un}}}\min_{\mathbf{F}_{\text{un}}}\sum_{e\in\mathcal{E}}&\mathbf{F}_{e}^{\intercal}(K_{e}(\mathbf{u}_{e},\mathbf{u}_{e})+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{F}_{e}+\log\det(K_{e}(\mathbf{u}_{e},\mathbf{u}_{e})+\sigma_{\epsilon}^{2}I)\\ &\text{s.t.}\qquad\delta_{0}^{\intercal}\mathbf{F}=0.\\ \end{split}$ (22)
@@
$\begin{split}F&=\widehat{K}^{-1}\widehat{D_{0}}\left(\widehat{D_{0}}^{\intercal}\widehat{K}^{-1}\widehat{D_{0}}\right)^{-1}b,\\ \lambda&=-(\widehat{D_{0}}^{\intercal}\widehat{K}^{-1}\widehat{D_{0}})^{-1}b.\\ \end{split}$ (31)
After reshaping $F$ [^3] That is, we reshape the final result to undo the vec($\cdot$) operation. and substituting back into the GP recovery problem in (22), we arrive at the following optimization problem:
@@ 4.3 Training details
We minimize the MLE-based objective in (32) with Adam [22], a first-order optimizer capable of handling nonconvex, nonlinear optimization problems. We use an initial learning rate of $10^{-3}$ and a StepLR scheduler that scales the learning rate by 0.98 every 10,000 epochs.
The trainable model parameters $\theta$ consist of edge-specific length scale parameters $\ell_{e}$ for all $e\in E,$ and a global noise variance $\sigma_{\epsilon}^{2}.$ To enforce positivity, we optimize the logarithms ~$\tilde{\ell}_{e}=\log{\ell_{e}}$ and $\tilde{\sigma}^{2}_{\epsilon}=\log{\sigma^{2}_{\epsilon}}$, and exponentiate them inside the computational graph. We initialize each $\ell_{e}$ to the median pairwise distance between training inputs on that edge and $\sigma_{\epsilon}^{2}$ to $\exp(-20)$, or approximately $92e-$ 9.
@@ 4.5 Error estimates
A core strength of our approach is the rigorous uncertainty quantification made possible by working with GPs. In 2, we establish formal bounds on the worst-case error between the minimizer of our optimization problem $\hat{f}$ and the true function $f$ governing the dynamics of a given edge. Following standard practice in (nonlinear) optimal recovery, we assume that $f$ lies in the RKHS space associated with our chosen kernel $K(\cdot,\cdot)$ [3, 14, 20, 24, 32, 40]. In other words, we assume that $f$ obeys the smoothness and structural properties encoded by $K(\cdot,\cdot),$ and can be approximated arbitrarily well in the RKHS norm by finite linear combinations of kernel sections $K(x_{i},\cdot).$
@@
$\lvert f(x)-\hat{f}(x)\rvert\leq\sigma(x)\left\lVert f\right\rVert_{\mathcal{H}_{K}}+\sigma_{n}\left\lVert K(x,\mathbf{X})[K(\mathbf{X},\mathbf{X})+\sigma_{\epsilon}^{2}I]^{-1}\right\rVert_{2}\sqrt{2\log(2/\delta)},$ (36)
@@
$\hat{f}(\cdot)$ $=K(\cdot,\mathbf{X})[K(\mathbf{X},\mathbf{X})+\sigma_{\epsilon}^{2}I]^{-1}\mathbf{Y}.$ (37)
@@
$e_{\text{det}}$ $=\langle f,K(x_{i},\cdot)\rangle_{\mathcal{H}_{K}}-\sum_{i=1}^{N}\phi_{i}(x)\langle f,K(x_{i},\cdot)\rangle_{\mathcal{H}_{K}},$ (45)
$=\langle f,K(x_{i},\cdot)\rangle_{\mathcal{H}_{K}}-\langle f,\sum_{i=1}^{N}\phi_{i}(x)K(x_{i},\cdot)\rangle_{\mathcal{H}_{K}},$ (46)
$=\langle f,K(x_{i},\cdot)\rangle_{\mathcal{H}_{K}}-\langle f,K(x,\mathbf{X})[K(\mathbf{X},\mathbf{X})+\sigma_{\epsilon}^{2}I]^{-1}K(\mathbf{X},\cdot)\rangle_{\mathcal{H}_{K}},$ (47)
@@
where (62) follows from the triangle inequality. To bound $\lvert e_{\text{noise}}(x)\rvert$, we use a Gaussian tail bound to address the unknown noise $\epsilon_{i}$. Recall that $e_{\text{noise}}(x)$ is a weighted sum of independent Gaussian random variables. Standard results for Gaussian concentration inequalities [39] tell us that:
@@
$\mathbb{P}\bigg{[}|e_{\text{noise}}|$ $\leq\sigma_{n}\left\lVert\phi(x)\right\rVert_{2}\sqrt{2\log(2/\delta)}\bigg{]}\geq 1-\delta.$ (64)
@@
$\lvert f(x)-\hat{f}(x)\rvert\leq\sigma(x)\left\lVert f\right\rVert_{\mathcal{H}_{K}}+\sigma_{n}\left\lVert\phi(x)\right\rVert_{2}\sqrt{2\log(2/\delta)}.$ (65)
@@
The analysis here does not account for uncertainty stemming from the inferred input to the GPs, and thus assumes that these inferred variables are accurate. Recall that the interior vertex values (and thus the inputs to the GP) are unknown and are estimated via the process described in Section 4.4. As a potential extension, this additional source of uncertainty could be incorporated by linearizing the problem around the maximum a posteriori (MAP) estimate, as in [7].
The error bounds in 2 make the common assumption that $f\in\mathcal{H}_{K}$. In practice, the true function $f$ might lie outside of the RKHS associated with our chosen kernel $K(\cdot,\cdot)$. In this case, often referred to as model misspecification, the induced RKHS should be interpreted as a “hypothesis space,” and our approach returns the function $\hat{f}\in\mathcal{H}_{K}$ closest to the projection of $f$ onto $\mathcal{H}_{K}$. For most commonly used kernels, $\mathcal{H}_{K}$ is dense in the space of continuous functions, and is thus rich enough to provide a suitable approximation of $f$ [36]. Furthermore, [14] show that many types of model misspecification either do not affect error estimate bounds at all, or only affect them to a minor extent.
@@ 5.1 Toy circuit
Figure 2: Histogram of $\log_{2}(\lvert\text{error}\rvert/\lvert\text{bound}\rvert)$ for different test-set sizes (color-coded). Values below zero indicate that the estimated bound safely over-predicts the true error; only a few cases in the sample of size 500 lie above zero, as expected from the $95\%$ confidence level.
Figure 2 presents the results of a simple scaling test that demonstrates the behavior of our approach as the test set increases in size. Specifically, we show a histogram of the ratio of actual test error to the estimated error bound. The ratios concentrate between $[-3,-7],$ meaning that the true error is typically between 8 and 128 times smaller than the predicted worst-case bound. Only when 500 test points are sampled do a handful exceed the bound, an outcome consistent with the nominal failure probability $\delta=0.05$. In other words, the estimated error bound is slightly conservative for most realistic test sizes but remains probabilistically valid.

### v2

@@ 1 Introduction
Multiphysics simulations are often composed of independent submodules exchanging information or fluxes through shared interfacial states [dvorak2005unifying]. A modular “systems-of-systems” perspective is therefore attractive for the prediction, design, and control of large-scale systems whose sheer complexity renders traditional simulation intractable. For example, the Energy Exascale Earth System Model (E3SM) forecasts planetary climate dynamics by coupling submodels representing the atmosphere, ocean, sea and land ice, and terrestrial systems. These submodels interact by exchanging mechanical and transport fluxes determined by their physical states [leung2020introduction]. Similarly, embedded electronic systems rely on individual circuit components that exchange voltage and current information, enabling the construction of integrated systems made up of billions of transistors.
@@
This exchange of state across subsystem boundaries can be formalized using Dirichlet-to-Neumann (D2N) maps. These maps arise in linear elliptic boundary value problems, where solving a partial differential equation (PDE) with given Dirichlet boundary conditions yields corresponding fluxes at those boundaries [natarajan1995domain]. D2N maps thus provide a natural abstraction for characterizing interactions between submodels sharing common interfaces. Many established domain decomposition techniques—including mortar methods [arbogast2007multiscale, bernardi1993domain], FETI methods [farhat2001feti], and hybridizable methods [cockburn2009unified]—can be viewed within the D2N framework. The rigorous theoretical foundations underpinning traditional D2N maps offer convergence and stability guarantees when coupling large legacy simulation codes [jiang2024structure, sockwell2020interface]. However, the evaluation of a D2N map traditionally requires computationally intensive solves of the governing equations within each subdomain, which quickly becomes prohibitively expensive in multi-physics and multiscale settings.
@@
Recent efforts have therefore turned to data-driven schemes that replace costly subdomain models with a learned approximation of the D2N map [bochev2024dynamic, hall2021ginns, jiang2024structure]. While such surrogates dramatically reduce computational cost, they generally lack the theoretical guarantees of traditional D2N methods. The absence of rigorous a-priori error or stability guarantees for data-driven D2N surrogates poses significant challenges for verification, validation, and reliable deployment of learned models.
@@
We address this gap by combining discrete exterior calculus (DEC) [TRASK2022110969] with computational graph completion (CGC) — a generalization of the optimal recovery problem (ORP) to the nonlinear graph setting [owhadi2022computational]. Our framework learns flux-state relationships on graphs while producing analytically and computationally tractable posterior distributions that characterize uncertainty in the D2N map. Together, CGC and DEC deliver an efficient optimization strategy along with closed-form expressions for prediction error and out-of-distribution inference, re-introducing rigorous guarantees into data-driven D2N coupling.
@@ 2 Optimal recovery on graphs via Gaussian processes
Our problem of learning a functional relationship from limited and potentially noisy data fits broadly within the framework of the optimal recovery problem (ORP). In mathematics, ORP involves finding the best approximation of an unknown function from limited observations. Given data generated by an unknown true function $f$, which we assume lies in some space $\mathcal{F}$ endowed with norm $\|\cdot\|$, classical ORP identifies the candidate $\hat{f}\in\mathcal{F}$ that minimizes the worst-case relative error:
@@
This topic is well-explored in the literature [donoho_statistical_1994, micchelli_survey_1977, Micchelli1976, Owhadi_Scovel_2019]. Variations of this problem include sensor reconstruction problems (e.g., [reconstruction]) and boundary value inverse problems (e.g., [ibvp_2000]). More recently, the computational graph completion (CGC) framework [owhadi2022computational] used Gaussian processes (GPs) to generalize the ORP to the graph setting, where constraints may be nonlinear.
GPs offer a powerful probabilistic framework for approaching the optimal recovery problem. A GP defines a prior distribution over a space of functions that can be updated with observed data to obtain a posterior distribution. The posterior mean provides the optimal recovery of the unknown function given the data, while the posterior covariance quantifies uncertainty. This makes GPs particularly useful for applications where both function approximation and uncertainty quantification are crucial. In this section, we provide a brief overview of Gaussian process regression, but we refer the interested reader to [rasmussen_gaussian_2008] for a more comprehensive but accessible treatment.
@@
[…]
@@
A GP is characterized by a covariance (kernel) function $K(\cdot,\cdot)$, which must be symmetric and positive-definite. According to the Moore–Aronszajn Theorem, each kernel $K(\cdot,\cdot)$ uniquely induces a corresponding reproducing kernel Hilbert space (RKHS) $\mathcal{H}_{K}$ [aronszajn_theory_1950]. This RKHS contains precisely the functions spanned by the GP kernel, and functions drawn from the GP prior belong to $\mathcal{H}_{K}$ with high probability. Working in $\mathcal{H}_{K}$ therefore yields solutions that are regularized according to the smoothness properties inherited from $K(\cdot,\cdot)$. Thus, prior information about the nature of the true function $f$ can be incorporated into the choice of kernel. The squared exponential RBF kernel:
@@
in the presence of noise. By the Representer Theorem, the solution to this problem lies in the span of kernel evaluations at the data points, dramatically simplifying the problem [scholkopf_generalized_2001]. Indeed, the GP posterior mean is exactly the minimizer of the RKHS-regularized problem, ensuring both smoothness and consistency with the data. In 1, we state the theorem that yields a closed-form solution for this minimization problem.
@@ 3 Discrete exterior calculus
DEC extends the framework of differential geometry and exterior calculus to discrete structures like graphs, providing a precise language through which we can define differential operators (e.g., gradient, divergence, and curl) in a discrete setting [desbrun_kanso_tong_2008, hirani_thesis_2003]. It is particularly well-suited for modeling flows on graphs, where quantities such as mass, heat, or electrical charge must obey conservation laws. In prior work, DEC has been used to learn physics from data [TRASK2022110969]. In our framework, we leverage DEC to constrain our graph-based GP recovery problem, ensuring that the learned functions obey conservation principles across the graphs we consider. This integration yields physically meaningful predictions while maintaining the flexibility of GPs.
@@
[…]
@@ 4.2 Formulation of the optimization problem
[…]
@@
[…]
@@
[…]
After reshaping FF [^3] That is, we reshape the final result to undo the vec($\cdot$) operation. and substituting back into the GP recovery problem in (22), we arrive at the following optimization problem:
@@ 4.3 Training details
We minimize the MLE-based objective in (32) with Adam [KingmaBa2014], a first-order optimizer capable of handling nonconvex, nonlinear optimization problems. We use an initial learning rate of $10^{-3}$ and a StepLR scheduler that scales the learning rate by 0.98 every 10,000 epochs.
The trainable model parameters $\theta$ consist of edge-specific length scale parameters $\ell_{e}$ for all $e\in E,$ and a global noise variance $\sigma_{\epsilon}^{2}.$ To enforce positivity, we optimize the logarithms $\tilde{\ell}_{e}=\log{\ell_{e}}$ and $\tilde{\sigma}^{2}_{\epsilon}=\log{\sigma^{2}_{\epsilon}}$, and exponentiate them inside the computational graph. We initialize each $\ell_{e}$ to the median pairwise distance between training inputs on that edge and $\sigma_{\epsilon}^{2}$ to $\exp(-20)$, or approximately 2e92e-9.
@@ 4.5 Error estimates
A core strength of our approach is the rigorous uncertainty quantification made possible by working with GPs. In 2, we establish formal bounds on the worst-case error between the minimizer of our optimization problem $\hat{f}$ and the true function $f$ governing the dynamics of a given edge. Following standard practice in (nonlinear) optimal recovery, we assume that $f$ lies in the RKHS space associated with our chosen kernel $K(\cdot,\cdot)$ [azarnavid_reproducing_2019, fiedler_practical_2023, kakade_worst-case_2005, leung2020introduction, reed_error_2024, zhou_gaussian_2023]. In other words, we assume that $f$ obeys the smoothness and structural properties encoded by $K(\cdot,\cdot),$ and can be approximated arbitrarily well in the RKHS norm by finite linear combinations of kernel sections $K(x_{i},\cdot).$
@@
$\lvert f(x)-\hat{f}(x)\rvert\leq\sigma(x)\left\lVert f\right\rVert_{\mathcal{H}_{K}}+\sigma_{\epsilon}\left\lVert K(x,\mathbf{X})[K(\mathbf{X},\mathbf{X})+\sigma_{\epsilon}^{2}I]^{-1}\right\rVert_{2}\sqrt{2\log(2/\delta)},$ (36)
@@
[…]
@@
$e_{\text{det}}$ $=\langle f,K(x,\cdot)\rangle_{\mathcal{H}_{K}}-\sum_{i=1}^{N}\phi_{i}(x)\langle f,K(x_{i},\cdot)\rangle_{\mathcal{H}_{K}},$ (45)
$=\langle f,K(x,\cdot)\rangle_{\mathcal{H}_{K}}-\langle f,\sum_{i=1}^{N}\phi_{i}(x)K(x_{i},\cdot)\rangle_{\mathcal{H}_{K}},$ (46)
$=\langle f,K(x,\cdot)\rangle_{\mathcal{H}_{K}}-\langle f,K(x,\mathbf{X})[K(\mathbf{X},\mathbf{X})+\sigma_{\epsilon}^{2}I]^{-1}K(\mathbf{X},\cdot)\rangle_{\mathcal{H}_{K}},$ (47)
@@
where (62) follows from the triangle inequality. To bound $\lvert e_{\text{noise}}(x)\rvert$, we use a Gaussian tail bound to address the unknown noise $\epsilon_{i}$. Recall that $e_{\text{noise}}(x)$ is a weighted sum of independent Gaussian random variables. Standard results for Gaussian concentration inequalities [vershynin2018] tell us that:
@@
$\mathbb{P}\bigg[|e_{\text{noise}}|$ $\leq\sigma_{\epsilon}\left\lVert\phi(x)\right\rVert_{2}\sqrt{2\log(2/\delta)}\bigg]\geq 1-\delta.$ (64)
@@
$\lvert f(x)-\hat{f}(x)\rvert\leq\sigma(x)\left\lVert f\right\rVert_{\mathcal{H}_{K}}+\sigma_{\epsilon}\left\lVert\phi(x)\right\rVert_{2}\sqrt{2\log(2/\delta)}.$ (65)
@@
The analysis here does not account for uncertainty stemming from the inferred input to the GPs, and thus assumes that these inferred variables are accurate. Recall that the interior vertex values (and thus the inputs to the GP) are unknown and are estimated via the process described in Section 4.4. As a potential extension, this additional source of uncertainty could be incorporated by linearizing the problem around the maximum a posteriori (MAP) estimate, as in [chen_gaussian_2024].
The error bounds in 2 make the common assumption that $f\in\mathcal{H}_{K}$. In practice, the true function $f$ might lie outside of the RKHS associated with our chosen kernel $K(\cdot,\cdot)$. In this case, often referred to as model misspecification, the induced RKHS should be interpreted as a “hypothesis space,” and our approach returns the function $\hat{f}\in\mathcal{H}_{K}$ closest to the projection of $f$ onto $\mathcal{H}_{K}$. For most commonly used kernels, $\mathcal{H}_{K}$ is dense in the space of continuous functions, and is thus rich enough to provide a suitable approximation of $f$ [steinwart]. Furthermore, [fiedler_practical_2023] show that many types of model misspecification either do not affect error estimate bounds at all, or only affect them to a minor extent.
@@ 5.1 Toy circuit
Figure 2: Histogram $\log_{2}(\lvert\text{error}\rvert/\lvert\text{bound}\rvert) fo$r different test-set sizes (color-coded). Values below zero indicate that the estimated bound safely over-predicts the true error; only a few cases in the sample of size 500 lie above zero, as expected from the $95\%$ confidence level.
Figure 2 presents the results of a simple scaling test that demonstrates the behavior of our approach as the test set increases in size. Specifically, we show a histogram of the ratio of actual test error to the estimated error bound. The ratios concentrate between [3,7],[-3,-7], meaning that the true error is typically between 8 and 128 times smaller than the predicted worst-case bound. Only when 500 test points are sampled do a handful exceed the bound, an outcome consistent with the nominal failure probability $\delta=0.05$. In other words, the estimated error bound is slightly conservative for most realistic test sizes but remains probabilistically valid.